### v2

# We Can Hide More Bits: The Unused Watermarking Capacity in Theory and in Practice

Abstract
Despite rapid progress in deep learning–based image watermarking, the capacity of current robust methods remains limited to the scale of only a few hundred bits. Such plateauing progress raises the question: How far are we from the fundamental limits of image watermarking? To this end, we present an analysis that establishes upper bounds on the message-carrying capacity of images under PSNR and linear robustness constraints. Our results indicate theoretical capacities are orders of magnitude larger than what current models achieve. Our experiments show this gap between theoretical and empirical performance persists, even in minimal, easily analysable setups. This suggests a fundamental problem. As proof that larger capacities are indeed possible, we train Chunky Seal, a scaled-up version of Video Seal, which increases capacity $4\times$ to ${1024}\,\text{bits}$, all while preserving image quality and robustness. These findings demonstrate modern methods have not yet saturated watermarking capacity, and that significant opportunities for architectural innovation and training strategies remain.
Figure 1: Existing image watermarking models have capacities well under what this paper suggests to be possible. Shown are theoretical bounds on watermarking capacity under a PSNR constraint alone (thick line) and in combination with robustness requirements (thin lines). Recent methods operate far below the achievable bounds, often by orders of magnitude, as seen in the log-scale inset. Our proposed Chunky Seal (1024 bits) pushes capacity higher than prior work, but is still very far from the theoretical limits, indicating a large potential for future development.
1 Introduction
Content provenance, the act of identifying the origin and history of media, is increasingly important as generative tools for creating and editing text, audio and video content are becoming ever more accessible and powerful. The most robust and reliable content provenance solutions, including those in the C2PA industry standard (C2PA), combine cryptographically secure metadata, fingerprinting, and invisible watermarking. Invisible watermarking is the key area of interest in both academia and industry. This paper focuses on the image domain, as it is the most mature, despite techniques also existing for text (Kirchenbauer et al., 2023; Dathathri et al., 2024) and audio (Chen et al., 2023a; San Roman et al., 2024). Image watermarking is already integrated into commercial models (e.g., Meta AI, Adobe Firefly and Google Imagen) and available for third-party use via open-source models (Zhu et al., 2018; Bui et al., 2023a) .
Invisible image watermarking embeds an imperceptible secret message of a certain capacity recoverable under a variety of perturbations, leading to an inherent capacity-quality-robustness trade-off. Classic methods used hand-crafted tools, such as the mid-frequencies of the discrete cosine transform (DCT) (Al-Haj, 2007; Navas et al., 2008). The advent of deep learning led to significant improvements in all three dimensions (Bui et al., 2023a; Luo et al., 2020; Tancik et al., 2020; Fernandez et al., 2024).
Yet, despite these techniques, it seems that progress has stagnated. State-of-the-art methods successfully embed around ${100{-}200}\,\text{bits}$ in a relatively imperceptible way (i.e., Peak Signal-to-Noise Ratio, PSNR, above ${40}\,\text{dB}$) while robust to perturbations. Improvements in quality and robustness continue, but they are only marginal, leading many to believe we are nearing the limits of what is possible.
Image watermarking indeed may indeed already be a solved problem. Unlike generative or discriminative models that can improve as data and parameters are scaled, watermarking has an inherent performance ceiling. Given an image resolution and a set of robustness constraints, there is a finite amount of information that can be embedded imperceptibly. The existence of this limit and the converging empirical performance of recent models naturally leads to a critical question:
Have we already reached the theoretical ceiling of watermarking performance?
To answer this question, we need to know what this limit actually is and to measure how close our models are to it. We address these challenges in the current paper and offer the following findings:
  1. i.
    We propose bounds on the capacity of watermarking under a PSNR constraint and robustness to linear augmentations, indicating capacities orders of magnitude larger than seen in practice.
  2. ii.
    Watermarking models are trained with constraints we cannot directly analyse so we retrain Video Seal (Fernandez et al., 2024), a state-of-the-art image and video watermarking model, to match our simplest theoretical setup: watermarking a single gray image under only a PSNR constraint. Yet, Video Seal fails to encode even ${1024}\,\text{bits}$, when we successfully encode ${2048}\,\text{bits}$ with a linear model, ${32{,}768}\,\text{bits}$ by tiling lower-resolution watermarks, and ${456{,}509}\,\text{bits}$ with a handcrafted model. This indicates severe structural limitations.
  3. iii.
    With the standard quality and robustness constraints, we train Chunky Seal, a simple scale-up of Video Seal, which embeds ${1024}\,\text{bits}$ while maintaining similar robustness and image quality.
Therefore, our theory and experiments show that
It is possible to achieve much higher capacities than we currently have,
although that might require innovation in architectures and training.
2 Related work
Classic principled methods. Early research on image watermarking was dominated by hand-crafted signal processing techniques, grounded in well-understood mathematical and perceptual models. These methods could operate directly in the pixel domain (Van Schyndel et al., 1994; Bas et al., 2002), or in transform domains, most commonly the discrete cosine transform (DCT, Bors and Pitas, 1996; Piva et al., 1997) and the discrete wavelet transform (DWT, Xia et al., 1998; Barni et al., 2001), as well as combinations of the two (Navas et al., 2008; Feng et al., 2010; Zear et al., 2018). A key insight was that perceptually significant frequencies tend to be preserved under transformations (Cox et al., 1997). Other schemes, such as (Ni et al., 2006), introduced perturbations to all pixels with specific values to enable very large payloads. Despite being principled and accompanied by theoretical guarantees, these methods have limited robustness to perturbations and, in some cases, cause noticeable image degradation.
Deep learning based-watermarking. With the development of deep learning techniques for computer vision, it was only natural to extend them to image watermarking. Vukotić et al. (2018) proposed adversarially attacking a fixed image extractors, an idea later built upon by Fernandez et al. (2022) and Kishore et al. (2022). A similar iterative approach, albeit lacking robustness, was attempted for steganography by Chen et al. (2023b). LISO operates by iteratively optimizing the embedding of information into images, but its performance degrades sharply under even mild compression. Notably, the only robustness-aware variant, LISO-JPEG, achieves merely $\sim$1 bpp at 19.7 dB PSNR, which is well below practical watermarking quality and is only robust to JPEG compression with CRF 80. This demonstrates that LISO is non-practical for real-world watermarking applications due to its weak robustness and low fidelity.
More popular and successful methods train purpose-built neural networks. Early convolutional models, such as those introduced by Mun et al. (2017) and Zhu et al. (2018), established the feasibility of CNN-based watermarking. Subsequently, architectures based on U-Net (Ronneberger et al., 2015) gained prominence, leveraging multi-resolution representations and residual connections to enable greater network depths. Recent work has also explored watermarking in the latent space of diffusion models Bui et al. (2023b). Advances in perceptual loss design have proven critical: careful loss selection and tuning improved both image quality (Bui et al., 2023a; Fernandez et al., 2024; Xu et al., 2025) and robustness (Tancik et al., 2020; Jia et al., 2021; Pan et al., 2024). Beyond robustness, methods for watermark localization and multi-message embedding have been proposed (Sander et al., 2025; Wang et al., 2024; Zhang et al., 2024), while add-on techniques such as adversarial training (Luo et al., 2020) and attention mechanisms (Zhang et al., 2020) further enhance performance. Although combining image generation with watermarking has also been studied (Fernandez et al., 2023; Wen et al., 2023; Kim et al., 2023; Hong et al., 2024; Ci et al., 2024), such generative approaches are beyond the scope of this work.
Information-theoretic capacity of watermarking
A natural foundation for analyzing watermarking capacity is the seminal framework of Gel’fand and Pinsker (1980), which characterizes communication over channels with random parameters. In the watermarking context, the cover image acts as the channel state: it is known to the encoder (the hider) but not to the decoder. The Gel’fand–Pinsker theorem expresses the capacity of such channels as
$C=\max_{p(u|s),\,x(u,s)}\bigl[I(U;Y)-I(U;S)\bigr],$
where $S$ denotes the cover signal, $U$ an auxiliary variable chosen by the encoder, $X$ the watermarked image, and $Y$ the attacked image. This framework underlies nearly all classical information-theoretic analyses of watermarking.
Early works such as Costa (1983) and Cohen and Lapidoth (2002) assume Gaussian cover signals and Gaussian, memoryless attacks. Under these assumptions, the decoder’s ignorance of the cover does not reduce capacity: the achievable rate depends only on watermark power and noise power. Quantization-based analyses (Chen and Wornell, 2002) operate under similarly stylized Gaussian models.
More sophisticated approaches were introduced by Moulin and O’Sullivan (2003) and Moulin and Koetter (2005), who cast watermarking as a strategic game between hider and attacker and defined the notion of hiding capacity. These formulations characterize the fundamental tradeoff between embedding distortion $D_{1}$ and attack distortion $D_{2}$, and have inspired a long line of follow-up work (e.g., Somekh-Baruch and Merhav, 2004 and Merhav, 2005). However, these analyses remain grounded in Euclidean or Hamming distortion metrics and in pixelwise probabilistic models of the cover and the attack. Even extensions that allow more complex or adversarial behavior—such as pixel permutations—still operate within a per-pixel, additive-noise paradigm.
These classical works provide valuable theoretical insight, but they share the same foundational assumption: watermarking is modeled as a channel coding problem over an idealized Euclidean signal space, with attacks represented as noise in that space. As discussed next, this assumption diverges sharply from the realities of digital images and modern watermark attacks.
The need for a new framework for watermarking capacity
The classical Gel’fand and Pinsker (1980) framework treats watermarking as a per-pixel communication problem: attacks are modeled as additive or memoryless noise, distortions are measured in Euclidean or Hamming metrics, and images are assumed to be continuous signals with well-defined probability densities. The capacity expression
I(U;Y)I(U;S)I(U;Y)-I(U;S)
reveals the central limitation of this view: mutual information depends on the absolute probability distribution of natural images. Such distributions are unknown, representation-dependent (RGB vs. YCbCr, gamma correction, quantization, compression), and fundamentally inestimable. Because $I(U;Y)$ and $I(U;S)$ require these true underlying densities—not empirical or relative statistics—Gel’fand–Pinsker capacities cannot be instantiated for real images.
Moreover, the dominant failure modes of watermarking systems are not pixelwise perturbations but geometric transformations: cropping, resampling, rotations, perspective changes, and non-rigid warps that move pixels rather than modify their intensities. These structured distortions introduce strong spatial dependencies that fall far outside memoryless noise models and cannot be meaningfully captured by $\ell_{2}$ or related Euclidean metrics. Real images also inhabit discrete, quantized spaces such as $\{0,\ldots,255\}^{3}$, making differential-entropy–based Gaussian capacity formulas inapplicable. As a result, classical information-theoretic capacity predictions are tied to an idealized Euclidean signal model that digital images simply do not satisfy.
Because real attacks are predominantly geometric and because natural-image densities are inaccessible, classical information-theoretic models cannot yield realistic or actionable capacity estimates. They cannot answer practical questions such as:
How many bits can be reliably embedded in a $128\times 128$ image while remaining recoverable after rotations, cropping, or other structured geometric distortions?
Our goal is to develop a capacity theory grounded in the geometry of images—one that predicts how many bits can be reliably stored under realistic transformations such as affine and non-rigid warps, and that provides bounds and constructions reflecting the true failure modes of modern watermarking systems.
3 Bounds on watermarking capacity
In this section we establish fundamental limits of image watermarking. As watermarking requires every message to be encoded as a distinct image, the problem can be formalized in a geometric framework: images are points on a high-dimensional grid, and capacity is determined by the number of unique such points that do not violate given constraints. Watermarked images must be close to the cover image to be imperceptible and, to remain robust, messages must be represented with sufficient redundancy. Both requirements reduce the number of admissible images available for encoding. We begin with establishing the absolute maximum information that can be represented in an image (Section˜3.1). We then add a PSNR constraint in Sections˜3.2 and 3.3. We incorporate robustness to transformations such as cropping, rescaling, rotation and JPEG in Section˜3.4. Finally, we consider the effect of the data distribution on capacity (Section˜3.5).
3.1 Absolute capacity of the image space
Watermarking embeds a message $m$ into an image $\bm{x}$. Since each message must correspond to a distinct encoded image, the number of unique messages, that is, the watermarking capacity is limited by the number of distinct images. An $l$-bit message requires at least $2^{l}$ such images. We represent an image as a vector of length cwhcwh, where $c$ is the number of channels, $w$ is the width and $h$ the height, with each element having $2^{k}$ discrete levels when using $k$-bit colour depth. The tuple $(c,w,h,k)$ defines an image format. The set of all possible images in this format is $\mathcal{I}=\{0,1,\dots,\rho\}^{cwh}$ with $\rho=2^{k}-1$, which can be thought of as a finite grid [^1] The set of valid images is a finite subset of a lattice in $\mathbb{R}^{cwh}$, i.e., $\mathcal{I}{=}[0,\rho]^{cwh}{\cap}\{\sum_{i=1}^{cwh}a_{i}\bm{e}_{i}\mid a_{i}{\in}\mathbb{Z}\}$ with $\bm{e}_{i}$ the $i$-th unit vector, but we use the term grid for readability since no deeper lattice theory is required. of integer points in $\mathbb{R}^{cwh}$. This immediately gives us a trivial upper bound on watermarking capacity: since each message must correspond to a distinct watermarked image, it is not possible to embed more messages than there are distinct images.
Bound 1 (Absolute capacity of the image).
The capacity of images in the format $(c,w,h,k)$ is ${\texttt{capacity}\text{[in bits]}}=\log_{2}|\mathcal{I}|=\log_{2}\big((2^{k})^{cwh}\big)=cwhk\;\;\text{bits}.$
Bound˜1 simply states that the maximum number of embeddable bits is the uncompressed size of the image in bits, i.e., 1.57 Mbits in a ${256}{\times}{256}\text{px}$ colour image. We next introduce imperceptibility and robustness constraints to measure their effect on capacity.
(a) Cube fully inside sphere
(low PSNR, Bound˜2)
(b) Sphere fully inside cube
(high PSNR, Bounds˜3 and 4)
(c) Non-trivial intersection
(medium PSNR, Bounds˜6 and 5)
(d) Sphere at corner, i.e.,
arbitrary image (Section˜3.3)
Figure 2: The box-ball configurations of the PSNR-only constraint. The cube $C_{\mathcal{I}}$ represents the set of all images and the sphere is the PSNR ball centred at the cover. Their intersection determines the set of feasible watermarked images, with the cardinality being the watermarking capacity. (a), (b) and (c) are the cases with the cover image $\bm{x}$ at the centre of the cube $C_{\mathcal{I}}$ (gray image, resulting in highest capacity, Section˜3.2). (d) is the case of the worst-case cover $\bm{x}$, i.e., at the corner of $C_{\mathcal{I}}$ (Section˜3.3).
3.2 Capacity under a PSNR constraint
In practice, watermarked images must remain visually close to the cover image, ideally imperceptibly so. This limits the set of admissible images for encoding messages. A standard way to quantify distortion is the peak signal-to-noise ratio (PSNR), measured in dB. Requiring a minimum PSNR $\tau$ between the cover $x$ and the watermarked image $\tilde{x}$ is equivalent to bounding their $\ell_{2}$ distance (see App. 7 for the full derivation):
${\text{PSNR}}(x,\tilde{x})\;\geq\;\tau\quad\Longleftrightarrow\quad\|x-\tilde{x}\|_{2}\;\leq\;\epsilon(\tau),\text{ with }\epsilon(\tau)=\rho\sqrt{cwh}\,10^{-\tau/20}.$ (1)
Interpreting PSNR as an $\ell_{2}$-ball constraint gives us an avenue for measuring the message-carrying capacity under it by considering the amount of integer points inside both the cube and this ball. Counting how many such points exist is not trivial, and we analyse the three possible cases (see Figure˜2): i. the ball is so large that it contains the entire cube (very low $\tau$); ii. the ball is small enough to lie fully inside the cube (high $\tau$); iii. the ball and cube partially overlap (medium $\tau$). We begin by assuming the cover image $\bm{x}$ lies at the centre of the admissible range, i.e., $\bm{x}_{g}=2^{k-1}\ \bm{1}$ as then the volume of the intersection (and thus the capacity) is maximized. In Section˜3.3 we will extend the analysis to arbitrary images.
3.2.1 Cube in ball (low PSNR)
When $\tau$ is low, $\epsilon(\tau)$ is large and the ball contains the entire cube $C_{\mathcal{I}}=[0,\rho]^{cwh}$. The PSNR constraint does not rule out any images, so the capacity is just the absolute maximum (Bound˜1):
Bound 2 (Gray image, PSNR constraint (low PSNR)).
The capacity of a gray image $\bm{x}_{g}$ under a very low minimum PSNR threshold $\tau$ is ${\texttt{capacity}\text{[in bits]}}=cwhk.$
Bound validity: When $\epsilon(\tau)\geq\nicefrac{{\rho}}{{2}}\sqrt{cwh}$, or equivalently when $\tau\leq 20\log_{10}2\approx{{6.02}\,\text{dB}}$.
3.2.2 Ball in cube (high PSNR)
When $\tau$ is high, the PSNR ball is fully inside the cube. The capacity is the number of integer points inside the ball. This amounts to counting the integer grid points in an $n$-ball of radius $\epsilon(\tau)$, a problem with no general closed form. In high dimensions cwhcwh and for sufficiently large radii $\epsilon(\tau)$, this is well approximated by the ball volume $\operatorname{Vol}B_{cwh}\left[\cdot,\epsilon(\tau)\right]$. See Section˜8 for details on the validity of this approximation.
Bound 3 (Gray image, PSNR constraint (high PSNR, volume approximation)).
The capacity of a gray image $\bm{x}_{g}$ under a high minimum PSNR threshold $\tau$ is approximately capacity[in bits] $\approx\log_{2}\operatorname{Vol}B_{cwh}\left[\cdot,\epsilon(\tau)\right]$ $=\log_{2}\frac{\pi^{cwh/2}\,\epsilon(\tau)^{cwh}}{\Gamma\!\left(\tfrac{cwh}{2}+1\right)}.$
[…]
For small radii the volume approximation becomes inaccurate. However, then there are relatively few integer points in the ball and we can explicitly count them, as long as the dimension cwhcwh of the ambient space is not too high. Instead of brute-force enumeration (which scales poorly), we use a method introduced by Mitchell (1966) leveraging symmetries for efficient counting (see Algorithm˜2).
Bound 4 (Gray image, PSNR constraint (high PSNR, exact count)).
For small $\epsilon(\tau)$ the capacity is ${\texttt{capacity}\text{[in bits]}}\;=\;\log_{2}\,\texttt{PointsInHypersphereMitchell}(\texttt{dim}=cwh,\texttt{radius}=\epsilon(\tau)).$
Bound validity: When $\epsilon(\tau)\leq\nicefrac{{\rho}}{{2}}$ (i.e., $\tau\geq 20\log_{10}(2\sqrt{cwh})$) and $\epsilon(\tau)$ small enough that exact counting is computationally feasible.
We use Bound˜4 whenever we can evaluate Algorithm˜2 in reasonable time, and otherwise Bound˜3. As shown in Figure˜3, the transition between the two regimes is smooth.
Figure 3: Discrepancy between the volume of a ball and the number of integer lattice points contained in it for small $\epsilon(\tau)$. When the radius is small, the volume-based approximation Bound˜3 underesimates the actual number of integer points. In such cases we use the exact count Bound˜4 instead. Evaluated for a $16\times 16\times 3=768$-dimensional ball.
Figure 4: Watermarking capacity under PSNR constraints for a $\times16$px 3-channel cover image without robustness requirements. We show the case where the cover lies at the centre of the pixel range (left, Sections˜3.1 and 3.2) and at the extreme corner where pixels are saturated (right, Section˜3.3). In both cases we plot the family of bounds we established ranging from the trivial total capacity (Bound˜1) to volume approximations (Bounds˜3 and 7), exact lattice counts (Bounds˜4 and 8), and numerical integration for partial overlap (Bounds˜5 and 9). In the right panel, the arbitrary cover image case lies below the centred gray image case by at most cwhcwh bits i.e., at most one bit per pixel (1 bpp). The discontinuity between Bounds˜7 and 8 is due to the volume approximation undercounting the integer points on the faces of the cube.
3.2.3 Non-trivial intersection (medium PSNR)
For intermediate PSNR values $\tau$, $B_{cwh}\left[\bm{x}_{g},\epsilon(\tau)\right]$ and $C_{\mathcal{I}}$ intersect non-trivially. We can approximate this count by the volume of the intersection, using the same volume-based method as in Bound˜3. One can use exact volume computation (see Bound˜5 in Section˜10), though this tends to be numerically unstable. In practice, a simpler upper bound approximates it well by taking the minimum of the two trivial cases:
Bound 6 (Gray image, PSNR constraint (medium PSNR, approximation)).
[…]
Bound validity: When $\nicefrac{{\rho}}{{2}}\leq\epsilon(\tau)\leq\nicefrac{{\rho}}{{2}}\sqrt{cwh}$, or equivalently $20\log_{10}2\leq\tau\leq 20\log_{10}(2\sqrt{cwh})$.
As shown in Fig. 4 left, this simple upper Bound˜6 closely tracks the exact Bound˜5. Thus Bound˜6 is the practical choice going forward, while Bound˜5 is provided in the appendix for completeness. Figure˜4 left illustrates all the bounds from this section for a ${16}{\times}{16}\text{px}$ image. At ${45}\,\text{dB}$ these bounds give us roughly ${2000}\,\text{bits}$ of capacity (more than ${2.5}\,\text{bpp}$): orders of magnitude more than the ${0.001}\,\text{bpp}$ we see in practice (Figure˜1). For the more reasonable ${256}{\times}{256}\text{px}$ watermarking resolution, that maps to roughly ${500{,}000}\,\text{bits}$.
3.3 From central gray image to arbitrary cover images
In Section˜3.2 we assumed the cover lies at the centre of the pixel range, thereby maximizing the volume of the intersection between the PSNR ball and the cube $C_{\mathcal{I}}$. Real images, however, may be anywhere in $C_{\mathcal{I}}$. Being at the corner of $C_{\mathcal{I}}$ minimizes overlap with the ball and thus provides a lower bound valid for any image. When $\epsilon$ is not too large, exactly $\nicefrac{{1}}{{2^{cwh}}}$ of the PSNR ball centred at a corner of $C_{\mathcal{I}}$ remains inside $C_{\mathcal{I}}$. Although this may seem drastic, the penalty is in fact modest: at most cwhcwh bits, i.e., one bit per pixel. In Section˜11 we provide the formal bounds for this corner setting. Bound˜7 adapts Bound˜3, the volume approximation when the ball is fully in the cube. Bound˜8 is the analogue of Bound˜4, i.e., exact counting for small $\epsilon(\tau)$. Bound˜9 parallels Bound˜5 for the case when numerical integration is needed. As shown in Figure˜4, the gap from the gray-only image bounds is at most ${1}\,\text{bpp}$, thus:
Watermarking with a PSNR constraint should allow for capacity upwards of 2bpp
and does not explain the low capacities we observe in practice.
3.4 Adding robustness constraints
In practice, watermarking must balance imperceptibility with robustness: the message should survive common processing, like compression, resizing, cropping, rotation, etc. We wish to quantify the capacity reduction due to the robustness constraints. We consider linear transformations, which encompass most transformations used in practice. We also develop LinJPEG, a linearized version of JPEG, allowing us to study the effects of compression in the same setting (see Section˜12.4 for the construction).
Take a linear transformation $M\in\mathbb{R}^{cwh\times cwh}$ that maps an image $\bm{x}$ to a transformed $M\bm{x}$ and a quantization operation Q (typically an element-wise rounding or floor operation) to map the pixel values of $M\bm{x}$ to the valid images $\mathcal{I}$. Hence, we have the final transformed image $\bm{x}^{\prime}=\texttt{Q}[M\bm{x}]$. We need to find the subset of the possible watermarked images under only the PSNR constraint that map to unique valid images after applying $M$ and Q to them. The main complication in this setup is that Q is non-linear.
Heuristic bounds. A simple approach is to take a volumetric approach akin to Bounds˜3, 7, 5 and 9. This gives rise to a set of relatively simple heuristic bounds. We factor in how $M$ changes the volume and account for directions compressed by the transformation which destroy capacity as different watermarked images get collapsed together. We also account for directions fully collapsed by $M$ when it is singular. Finally, the stretched directions might result in some watermarked images being outside $C_{\mathcal{I}}$ after the transformation, leading to them being clipped. Bounds˜10, 11 and 12 use a heuristic based on the singular values of $M$ to account for the effect on capacity. Refer to Section˜12.2 for details on their derivation and properties. In Figure˜6 we plot these bounds for robustness to rotation, cropping and LinJPEG, showing that even under the most aggressive cropping, we should expect around ${0.5}\,\text{bpp}$ or almost ${100{,}000}\,\text{bits}$ for ${256}{\times}{256}\text{px}$ images.
Table 1: Conservative capacity bounds under robustness constraints for PSNR 42 dB. These values are calculated via Bound˜13 and are strongly conservative lower bounds on the capacity that is achievable while maintaining robustness to the respective transformations and PSNR under ${42}\,\text{dB}$.
Conservative capacity
Augmentation bpp for ${16}{\times}{16}\text{px}$ for ${256}{\times}{256}\text{px}$
Horizontal Flip 3.064 2,352 bits 602,353 bits
Crop&Rescale 25% 0.107 81 bits 20,958 bits
Crop&Rescale 50% 0.015 11 bits 3,013 bits
Crop&Rescale 75% 0.005 3 bits 904 bits
LinJPEG q=8 0.134 102 bits 26,273 bits
LinJPEG q=10 0.136 104 bits 26,757 bits
LinJPEG q=15 0.137 105 bits 27,020 bits
Rotation 15deg 0.084 64 bits 16,502 bits
Rotation 30deg 0.075 57 bits 14,676 bits
Rotation 45deg 0.083 64 bits 16,401 bits
Figure 5: Robustness constraints considered in this paper. The figure illustrates the linear transformations we evaluate: Horizontal Flip, Crop&Rescale, Rotation and Linearlized JPEG compression. We show them both at full resolution and at ${16}{\times}{16}\text{px}$ at which we compute the bounds in Section˜3.4.
Figure 6: Impact of robustness constraints on watermarking capacity. Capacity (in bits per pixel) under a PSNR constraint is shown for three families of transformations: Crop&Rescale (left), Rotation (center), and Linearized JPEG (right), using the heuristic bounds (Bounds˜12, 11 and 10). The red lines show the PSNR-only capacity bounds without robustness constraints Bounds˜2, 3, 5 and 4. Each transformation reduces capacity in proportion to its severity: smaller crop scales, larger rotations, or lower JPEG quality factors. Across all cases, robustness constraints reduce but do not eliminate the large theoretical capacity gap with current watermarking methods.
Conservative bounds. We can show cases where these heuristic bounds under-approximate and cases where they over-approximate the true capacity, e.g., Figures˜9 and 10. Thus, the true capacity under linear transformation could be much lower than these bounds predict. To ensure that this is not the case, we develop an actual lower bound: Bound˜13. While we reserve the details for Section˜12.3, this bound is based on over-approximating the set of images that can be quantized by Q to the same image after $M$ is applied to them. As a result, Bound˜13 is extremely conservative and unrealistic. We believe that despite Bounds˜10, 11 and 12 not being valid lower bounds, they are much closer to the true capacity. Still, we report the conservative bound in Table˜1: the most aggressive crop still leaves at least ${904}\,\text{bits}$ for ${256}{\times}{256}\text{px}$ images. For the other augmentations, the conservative capacity is much higher. Therefore,
Robustness to geometric transformations and compression significantly reduces the capacity
but cannot fully explain the low watermarking capacity of current models.
3.5 From single cover images to datasets and data distributions
In a blind watermarking setup, the decoder must operate without access to the original cover image, creating potential collisions: if multiple natural images (i.e., potential covers) are very close to each other in pixel space, a watermarked version of one cover could be identical to a watermarked version of another. To prevent such ambiguity, the total set of watermarked images within a given region (like the PSNR ball) must be partitioned among all the potential covers it contains. If there are $N$ possible covers, the capacity for each is reduced by $\log_{2}(N)$ bits. We estimate $N$ using neural compression models like VQ-VAE (Van Den Oord et al., 2017) and VQGAN (Esser et al., 2021), which upper-bound the number of perceptually distinct images. For instance, a ${256}{\times}{256}\text{px}$ image can be compressed into a $32{\times}32$ latent with a 1024-entry codebook (Muckley et al., 2023). This representation can express at most $1024^{32\times 32}=2^{10240}$ distinct images. Conservatively assuming all could fall in the PSNR ball of the considered image, capacity is reduced by $10,240$ bits, or about ${0.05}\,\text{bpp}$, on top of the ${1}\,\text{bpp}$ loss from Section˜3.3. Thus, from this perspective,
The data distribution has only a negligible effect on watermarking capacity
and cannot explain the low performance of current models.
This aligns with prior findings for Gaussian channels that decoder knowledge of the cover does not affect capacity (Costa, 1983; Chen and Wornell, 2002; Moulin and O’Sullivan, 2003).
4 Empirical performance is much lower than predicted
Section˜3 showed that capacities of over ${2}\,\text{bpp}$ at PSNR of ${40}\,\text{dB}$ without robustness constraints, and of ${0.5}\,\text{bpp}$ with robustness, are possible. Even under the very conservative Bound˜13 we still would expect capacities of at least ${0.01}\,\text{bpp}$. However, in practice, the models reported in the literature have significantly lower capacities (less than ${0.001}\,\text{bpp}$, Figure˜1). To understand the cause of this gap, this section asks:
Are existing models significantly under-performing relative to what is possible in practice,
or are our bounds too unrealistic?
There are five possible explanations of the large discrepancy between the performance we see in practice (Figure˜1) and the bounds in Section˜3:
  1. A.
    Real models might be near-optimal if we consider advanced robustness constraints. Real models are trained with combinations of more complex robustness constraints than what our theory considers. Possibly, if we could factor these additional robustness constraints, our bounds would also show much lower possible capacity.
  2. B.
    Real models might be near-optimal if we consider advanced perceptual constraints. Real models satisfy advanced perceptual constraints beyond just PSNR, e.g., SSIM (Wang et al., 2004) or LPIPS (Zhang et al., 2018). Possibly, if we could factor these perceptual constraints, our bounds would also show much lower possible capacity.
  3. C.
    Real models might be near-optimal if we consider real-world image distributions. Real models are trained and evaluated on real-world image distributions, not a single fixed image. We argued in Section˜3.5 why that should not be a problem, but perhaps we are wrong and the need for disambiguating the cover image reduces the watermarking capacity much more than expected.
  4. D.
    Our bounds overestimate capacity and cannot be approached empirically.
  5. E.
    We can do much better and push the Pareto front well beyond the current state-of-the-art. The architecture and training setups, or compute, we use in practice are not good enough to fully utilise the watermarking capacity of the images, even when factoring in for advanced robustness, perceptual and data distribution constraints.
To understand the cause of the gap between theoretical and real-world performance in image watermarking, we need to find out which of these hypotheses is the underlying cause. If it is A., B., C., D., or a combination of them, then it is possible that, indeed, the best current models are close to what is ultimately possible and we can expect only marginal further improvements. On the other hand, if the cause is E., then that means that there is plenty of space for significant improvements.
Figure 7: Video Seal fails to learn how to embed ${1024}\,\text{bits}$ into a gray image with only PSNR constraint, whereas a linear embedder and extractor learn how to embed ${2048}\,\text{bits}$. (Left.) Video Seal trained on a single solid gray image with only the detector and MSE losses. It learns to embed up to ${512}\,\text{bits}$ but fails to embed ${1024}\,\text{bits}$. (Centre.) Same setup as the left plot but trained on a reduced ${32}{\times}{32}\text{px}$ resolution. The performance is similar; hence, the Video Seal architecture fails to make use of the full resolution. (Right.) Replacing the embedder and decoder of Video Seal with a single linear layer each achieves 100% bit accuracy and PSNR above ${40}\,\text{dB}$ for both 1024 and ${2048}\,\text{bits}$: demonstrating that Video Seal indeed has structural limitations. The results for the sweeps over the learning rate and $\lambda_{i}$ are shown, with the best models highlighted.
Table 2: Video Seal fails to learn how to embed 1024 bits into a gray image with only PSNR constraint while a linear embedder and extractor learn how to embed 2048 bits. Numerical results for the best-performing runs from Figure˜7 and their respective hyperparameters, as well as the handcrafted embedder/decoder from Equation˜2 at four representative PSNR values.
Message size
Message size if
tiled to 256x256px
PSNR Bit acc. $\lambda_{i}$ lr
VideoSeal (256x256px, 600 epochs) 128 bits 53.45 dB 100.00% 0.5 5e-4
256 bits 53.98 dB 100.00% 1.0 5e-4
512 bits 51.45 dB 100.00% 0.5 5e-4
1024 bits 40.10 dB 89.63% 1.0 5e-5
VideoSeal (32x32px, 600 epochs) 128 bits 8192 bits 51.02 dB 100.00% 1.0 5e-4
256 bits 16384 bits 48.98 dB 100.00% 1.0 5e-4
512 bits 32768 bits 41.66 dB 100.00% 1.0 5e-5
1024 bits 65536 bits 29.66 dB 84.39% 0.1 5e-5
1024 bits 65536 bits 33.20 dB 83.86% 0.5 5e-5
1024 bits 65536 bits 34.63 dB 83.78% 1.0 5e-5
1024 bits 65536 bits 50.83 dB 50.60% 0.5 5e-4
Linear (256x256px, 50 epochs) 1024 bits 44.28 dB 100.00% 20.0 5e-4
2048 bits 40.40 dB 100.00% 12.0 5e-4
Handcrafted 623232 bits 36.00 dB 100.00%
551948 bits 38.00 dB 100.00%
456509 bits 42.00 dB 100.00%
311616 bits 48.00 dB 100.00%
4.1 The real-world complexity does not explain the performance gap
Let’s first address cases A., B., C., i.e., that our bounds cannot capture the complexity of the robustness, quality and data constraint with which real models are trained. While we cannot bring the real-world complexity to our analytical bounds, we can bring the models to the simplified theoretical setup.
More concretely, we take the simplest of setups: a single gray image with a PSNR constraint, as in Section˜3.2. We will use Video Seal as the base for our experiments (Fernandez et al., 2024), originally introduced as an image watermarking model with frame copying that generalizes to video. It was first demonstrated with a 96-bit capacity and was recently extended to a 256-bit open-source version, which we use as the strongest available baseline. To match the setup of Section˜3.2, we replace the dataset with a single solid gray image, remove all perceptual constraints but the MSE loss and remove all augmentations. We first retrain it for $n_{\text{bits}}=$ 128, 256, 512, and ${1024}\,\text{bits}$. We have hereby reduced the task to simply find a way to encode $n_{\text{bits}}$ into a single fixed image. From Figure˜4 we expect capacities of around ${600{,}000}\,\text{bits}$ at ${40}\,\text{dB}$ in this setup. Thus, the model should easily learn these much lower $n_{\text{bits}}$. We train with AdamW (Loshchilov and Hutter, 2019) with batch size 256 for 600 epochs, 1000 batches per epoch, cosine learning rate schedule with a 20-epoch warm-up, similarly to Video Seal. We sweep over the learning rate (5e-4, 5e-5, 5e-6) and $\lambda_{i}$, the MSE loss weight (0.1, 0.5, 1.0), with LR=5e-5 and $\lambda_{i}=0.5$ being the values used for training Video Seal.
The results of training Video Seal on a single gray image can be seen in Figure˜7 left and Table˜2. There are runs for the 128, 256 and 512 bit models that do achieve 100% bit accuracy and PSNR values above 42dB. However, Video Seal cannot even get to ${1024}\,\text{bits}$, far from what we expect from the bounds. This is surprising: the model cannot approach the theoretical bounds even after removing the complexities that supposedly make watermarking difficult. This means that neither A., B. nor C. can explain why we see such a gap between the theoretical and real-world performance.
4.2 Our simplest bounds are achievable, yet models struggle to get near them
Section˜4.1 showed that Video Seal cannot match the capacity predicted by the bounds in Section˜3.2 even when trained only on a single gray image and with no augmentations. Thus, the complexity of real world watermarking cannot explain the gap between the theoretical and real-world performance. This leaves us with two options: D. our bounds are wrong and unachievable, or E. our models are under-performing. There are a couple simple experiments that can demonstrate that we can get much closer to the bounds in Section˜3.2 and hence D. also does not explain the gap.
Linear embedder and extractor. We trained a simple linear embedder and extractor. The embedder gets the 1024 bit message (shifted and scaled to 1-1 and +1+1 values) and produces a $256{\times}256{\times}3$ watermark residual which gets added to the original gray image. Similarly, the decoder is a linear layer from the flattened $256{\times}256{\times}3$ image to 1024 outputs, which are thresholded to recover the message. We train only for 50 epochs, with the same learning rate values and $\lambda_{i}\in\{4,8,12,20\}$.
The results in Figure˜7 right and Table˜2 show the linear layer learns what Video Seal could not: 100% bit accuracy for ${1024}\,\text{bits}$ with PSNR of ${44}\,\text{dB}$. We also trained a linear model for ${2048}\,\text{bits}$ which achieved 100% bit accuracy. This shows that capacities beyond 512 bits are possible in practice (at least for a gray image and no robustness) and are learnable via gradient descent. All one needs is the right architecture.
Lower-resolution training and tiling. Our experiments reveal that Video Seal does not exploit the additional degrees of freedom available at higher image resolutions. When trained at ${256}{\times}{256}\text{px}$, the model achieves essentially the same capacity and PSNR as when trained at ${32}{\times}{32}\text{px}$ (see Section˜4.1 and Table˜2). To verify this, we train Video Seal in the setup of Section˜4.1 at ${32}{\times}{32}\text{px}$ using the same learning-rate and $\lambda_{i}$ sweeps for 600 epochs. As shown in Figure˜7 (centre) and Table˜2, the performance at ${32}{\times}{32}\text{px}$ is nearly identical to that at ${256}{\times}{256}\text{px}$: the 512-bit model reaches 100% bit accuracy with ${41.7}\,\text{dB}$, despite operating on $64\times$ fewer pixels. In other words, the effective capacity we observe at ${256}{\times}{256}\text{px}$ is comparable to what one would expect around ${20}{\times}{20}\text{px}$, confirming that the architecture fails to utilize the available resolution.
Because this setup does not require robustness to geometric or valuemetric transformations and we consider only gray images, we can use the ${32}{\times}{32}\text{px}$ model to demonstrate that higher capacities are possible. A simple tiling strategy suffices: each tile is embedded with an independent secret using the same model. The decoder similarly is applied per patch with the individual decoded messages concatenated to obtain the final combined message. Using ${256}{\times}{256}\text{px}$ as the reference size, tiling yields $64\times$ the capacity of the base ${32}{\times}{32}\text{px}$ model. Thus, tiling the 512-bit model—which already achieves 100% accuracy at ${41.7}\,\text{dB}$—produces a watermark with ${32{,}768}\,\text{bits}$ total capacity while maintaining the same PSNR (which is resolution-independent). This effective capacity of ${32{,}768}\,\text{bits}$ is already much closer to our bound of roughly ${600{,}000}\,\text{bits}$, though still only about ${0.167}\,\text{bpp}$.
It is interesting that the model could not learn at the ${256}{\times}{256}\text{px}$ resolution even for 1024 bits when it is clear that it is possible to embed 32,768 bits as seen here. More importantly, this shows that our bounds are not that far off and capacities of at least 32,768 bits are indeed possible.
Figure 8: Simple models outperform Video Seal on a gray image with only a PSNR constraint. Experiments from Section˜4 compare our theoretical bounds (Section˜3.2) against trained models. Video Seal falls well below the predictions, while a linear model performs slightly better and a tiled ${32}{\times}{32}\text{px}$ Video Seal is even better. Our handcrafted model nearly matches the bound.
Handcrafted embedder and extractor. We can do even better by manually crafting an embedder and extractor. The key observation is that mapping a hypercube to binary messages is easy. Take the ball of radius $\epsilon(\tau)=\rho\sqrt{cwh}\ 10^{-\tau/20}$ from Equation˜1. The half-side of the largest cube that can fit in this ball is $d=\epsilon(\tau)/\sqrt{cwh}=\rho\ 10^{-\tau/20}$. We have that each edge of the box contains a cwhcwh-dimensional grid of $q=2\lfloor d\rfloor+1=2\lfloor 2^{k}\ 10^{-\tau/20}\rfloor+1$ points per side. Hence, that gives us total capacity in bits of
$\log_{2}\left[\left(2\left\lfloor 2^{k}\ 10^{-\tau/20}\right\rfloor+1\right)^{cwh}\right]=cwh\log_{2}\left[2\left\lfloor 2^{k}\ 10^{-\tau/20}\right\rfloor+1\right]=cwh\log_{2}q,$ (2)
or $\log_{2}q$ bits per pixel. See Figure˜8 for a plot of that for different PSNR values. For ${42}\,\text{dB}$, and images of ${256}{\times}{256}\text{px}$ that gives us a capacity of ${456{,}509}\,\text{bits}$ (see Table˜2)  almost $14\times$ what we could embed with the ${32}{\times}{32}\text{px}$ tiling approach. Moreover, it gets us close to the theoretical bound. This scheme is easy to encode and decode: convert the binary string to $q$-nary representation and then map each digit in this representation to one of the axes of the box. The decoding is just as easy: get each digit by measuring where the watermark stands along each dimension, combine into a $q$-nary representation and convert back to binary.
Therefore, we can get much closer to the boundary, at least in the solid gray image case with PSNR constraint and no robustness requirements. Thus, case D., that our bounds are wrong and impossible to achieve, is unlikely. This leaves us with one possible explanation as to why models in practice do not exhibit performance anywhere near what our theory predicts. That would be option E:
Our models are likely significantly underperforming relative to what is possible in practice.
We likely can do much better and push the Pareto front well beyond the current state-of-the-art.
The architectures we use might have the wrong inductive biases or poor hyperparameter choices, our datasets might not be big enough, we might not be training long enough, or our models might be too small. Whatever the specific cause, image watermarking models clearly have significant room for performance improvement.
5 Better performance in practice is possible: Chunky Seal
While it remains possible that current models approach a theoretical limit under robustness and quality constraints, training a watermarking model with comparable quality and robustness but with substantially higher capacity would decisively rule this out. We take Video Seal (Fernandez et al., 2024) as the base model and train it for ${1024}\,\text{bits}$. We increased the embedding dimension to 2048, the U-Net channel multipliers from $[1,2,4,8]$ to $[4,8,16,32]$, and enabled watermarking in all three channels, not just the luma (Y) channel. This results in an embedder $90\times$ larger than the original Video Seal embedder. The ConvNeXt (Liu et al., 2022) extractor was similarly scaled: we increased the depths for each stage from $[3,3,9,3]$ (as in ConvNeXt-tiny) to $[3,3,27,3]$ (as in ConvNeXt-base), with their dimensions increased from $[96,192,384,768]$ to $[256,512,1024,2048]$. The stride of the first layer was reduced from 4 to 2. This results in an extractor that is $23\times$ larger than the original Video Seal extractor. Due to its significantly increased size, we name this model Chunky Seal. We train it at the original ${256}{\times}{256}\text{px}$ resolution. We apply gradient clipping with a maximum norm of 0.01, which proved critical for stabilizing training.
As shown in Table˜3, Chunky Seal shows image quality and robustness comparable to Video Seal across a wide range of distortions, while providing a $\bm{4}\bm{\times}$ higher message capacity (1024 vs. ${256}\,\text{bits}$). Despite its much larger capacity, Chunky Seal maintains nearly identical image quality across all metrics, and only slightly higher LPIPS. The robustness results further confirm that Chunky Seal sustains high bit-accuracy across transformations such as rotation, resizing, cropping, brightness and contrast changes, JPEG compression, and blurring, closely matching Video Seal. We emphasize that these results were achieved without hyperparameter tuning, whereas Video Seal was extensively optimized for quality and robustness.
Achieving $\bm{4\times}$ the capacity per pixel with comparable robustness and quality through simple scaling strongly suggests that substantially higher capacities are within reach using improved architectures and training strategies.
Table 3: Chunky Seal performance on images from SA-1B (Kirillov et al., 2023) at their original resolution. Chunky Seal has much higher capacity (${1024}\,\text{bits}$) than Video Seal while preserving its image quality and robustness on a wide variety of transformations. The improvement is driven by scaling the model size and its training. Extended results on SA-1B (Kirillov et al., 2023) and COCO (Lin et al., 2014) as well as qualitative results, are reported in Section˜15.1
Chunky Seal (ours) Video Seal 256bits
Capacity 1024 bits 256 bits
0.0052 bpp 0.0013 bpp
Embedder size 1022.7M 11.0M
Extractor size 773.7M 33.0M
[…]
6 Discussion
Given the theoretical and empirical results outlined in the previous sections and the nuanced interpretations, we would like to discuss some of the implications and the limitations of our findings.
Why maximise capacity? One could argue that we don’t need kilobytes of watermarking capacity; the roughly ${256}\,\text{bits}$ we currently have is plenty. However, higher capacities can open up new avenues for content provenance. For example, rather than using watermarks to encode a hash of a C2PA manifest so that we can recover it from a trusted third-party database (Collomosse and Parsons, 2024), we can instead encode the whole manifest itself, removing the need for registries.
Still, the implications of this paper are not limited to increasing the capacity of watermarking. Depending on the application, one might have a fixed length payload and robustness constraints and might want to maximise image quality; or fixed capacity and image quality and wanting to maximise robustness. We did not explicitly look at these cases but the inherent trade-offs between capacity, robustness and quality imply that if we succeed at significantly improving one of these axes, we can trade that improvement for the other two. In other words, by showing that we should be able to achieve orders of magnitude higher capacities, keeping the image quality and robustness fixed, it is not inconceivable that one would then be able to significantly increase image quality and/or robustness while maintaining capacity.
Implications on the nature of image data. In this work, we argue that the intrinsic capacity for embedding messages in images is significantly higher than current methods suggest. This proposition has two important implications. First, it confirms that the intrinsic dimensionality of the natural image data manifold is, in fact, low. Second, it suggests that neural compression techniques could achieve much higher compression rates than what is currently possible. Furthermore, our finding that watermarking capacity is an order of magnitude larger than even the highest-capacity state-of-the-art models helps explain the feasibility of applying and detecting multiple watermarks on a single image, as demonstrated in (Petrov et al., 2025).
Why doesn’t deep learning get us anywhere close to the bounds? Despite Chunky Seal obtaining much higher capacities than previous watermarking models, it nevertheless is still very far from the bounds for even the most aggressive augmentations, as illustrated in Figure˜1. This raises the question of why is it so difficult to get close to the bounds. From the controlled setup experiments in Section˜4 we saw that the causes are likely not the data distribution (Video Seal on a single gray image did not get anywhere close to the bounds), the resolution (Video Seal performed similarly at lower resolutions), or the augmentations (Video Seal performed similarly when augmentations were removed). That left the model architecture or other aspects of the training as unsuitable. Chunky Seal showed that scaling the model size might be part of the solution, but even that had only marginal success. The fact that simple linear embedder and extractors outperform Video Seal in the simple setup of Section˜4.2 indicates that we do have an architecture problem. This is not surprising as the embedder-extractors pair is effectively learning an identity map, something notoriously difficult for neural networks (He et al., 2016; Hardt and Ma, 2017). Therefore, we need more radical innovation in architectures, losses and training pipelines.
Towards principled watermark development, testing and benchmarking. Section˜4 illustrated some unexpected phenomena in Video Seal which likely also exist in other models. Namely, increasing the watermarking resolution would not enable higher capacity, nor would removing the augmentations. Hence, we propose a set of sanity checks that we believe a new generation of principled watermarking methods should pass.
When trained on solid gray images with no augmentations, a principled watermarking method should have its capacity in bits increasing linearly with the number of pixels, should decrease linearly with increasing PSNR and should reach at least the performance of the linear and the handcrafted baselines from Section˜4. When trained on gray images with a single augmentation, its capacity should decrease as predicted with stronger augmentations, e.g., cropping to $\nicefrac{{1}}{{4}}$ of the area should lead to a 4-fold reduction in capacity and LinJPEG with quality 15 should lead to a 2-fold reduction. These properties are not a guarantee for good performance of watermarking in practice but necessary conditions to be (close to) Pareto-optimal. We hope that by aiming to satisfy these properties, we, as a community, will produce new generation of watermarking models with orders of magnitude higher capacity and/or significantly better image quality.
Limitations. In spite of our attempt to be as comprehensive as possible, a few limitations remain in the analysis and experiments in the present paper. For one, we only consider image watermarking. While our results likely extend to video (which might have even larger capacities due to the temporal dimension), it is not clear what the bounds on audio or text watermarking would be. Our theoretical bounds in Section˜3 were limited to the setups we could study analytically and thus miss many aspects used in practice, such as combinations of transformations, advanced image quality techniques and accurate assessment of the effect of data distribution and the model size. Therefore, this leaves room for more advanced theoretical bounds that can be developed and that would be of significant value. Some of our bounds require numerical integration in very high dimensions and hence cannot be accurately evaluated at higher resolutions. While we resorted to using bits per pixel values computed at the manageable ${16}{\times}{16}\text{px}$ resolution, better numerical methods could be beneficial. Our main robustness bounds are also heuristics rather than formal upper bounds. While we do also have valid upper bounds they are extremely conservative and unrealistic, leaving plenty of space for improvement. Finally, while Chunky Seal does outperform Video Seal, it is also much larger and induces significant latencies which make it difficult to deploy at scale or on-device. Therefore, we need to achieve better watermarking performance, ideally without needing significant scaling up of our models.
Acknowledgments
This work would not have been possible without the feedback and support of Alex Mourachko, Sylvestre-Alvise Rebuffi, Tuan Tran and Valeriu Lacatusu. We would like to especially thank Yoshinori Aono, for answering all our questions and assisting with implementation of some of his results. We thank Matthew Muckley and Karen Ullrich for early discussions about this project. The authors are also grateful to Teddy Furon for pointing us to some relevant literature. AP acknowledges partial support by the EPSRC Centre for Doctoral Training in Autonomous Intelligent Machines and Systems (EP/S024050/1).